OpenStudy (anonymous):

What is the missing number? Five-sixths is equal to how many twenty-fourths? 4 5 8 twenty

OpenStudy (nincompoop):

\[\frac{ 5 }{ 6 } = \frac{ x }{ 24 }\]

OpenStudy (anonymous):

multiply each by 4. 5/6 = x/24 6 x 4 = 24. 5 x 4 = 20. so your answer is 20

OpenStudy (anonymous):

20